OpenStudy (anonymous):

If y varies inversely as x, find the constant of variation if y = 36 when x = 9. A. 27 B. 14 C. 324 D. 4

OpenStudy (whpalmer4):

If \(y\) varies inversely as \(x\) then we know that \(y = k/x\) where \(k\) is the constant of variation. You know a pair of values for \(x, y\) so use them to find \(k\).

OpenStudy (anonymous):

so 324 ?

OpenStudy (whpalmer4):

Let's check:\[y=\frac{k}{x}\]\[36=\frac{k}{9}\]\[36*9=k = 324\checkmark\]

OpenStudy (anonymous):

easy enough, thank you! i just didn't remember the equation right, thanks again (:

OpenStudy (anonymous):

\[\bf y=\frac{k}{x} \rightarrow 36=\frac{k}{9} \implies k=?\]